OpenStudy (anonymous):

What is the derivative of 1/(squareroot of x) and why?

OpenStudy (bahrom7893):

What is: \[\frac{1}{\sqrt{x}}\] In terms of powers?

OpenStudy (bahrom7893):

What I mean is, rewrite it as \[x^a\]

OpenStudy (anonymous):

x^(-1/2). Is that the derivative though? I remember my professor telling me that it's something like 1/((2)(square root of x)). Maybe I'm mistaken

OpenStudy (bahrom7893):

Well it is x^(-1/2), now just use the power rule.

OpenStudy (jhannybean):

\[\large \frac{d}{dx} x^n = \frac{nx^{n-1}}{n-1}\]

OpenStudy (bahrom7893):

uhm get rid of the denominator..

OpenStudy (bahrom7893):

\[n*x^{n-1}\]

OpenStudy (anonymous):

PLEASE HELP AsAP 1. a statement that you prove true is a(n) ? 2. When a conditional and it's converse are true, you can write them as a single true statement called a(n) ?

OpenStudy (jhannybean):

.... yes.